OpenStudy (anonymous):

Does this answer look right?? 5 ___ 8-✔5 My answer: 5✔5 ___ 3 a minute ago

OpenStudy (lgbasallote):

\[\frac 5{8 - \sqrt 5}\] if you multiply by conjugate \[\frac 5{8 - \sqrt 5} \times \frac{8+\sqrt 5}{8+\sqrt 5}\] is that what you did?

OpenStudy (anonymous):

Oh I just multiplied by the ✔5... So u do the whole thing to top n bottom?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

if you have sum/difference in the denominator you multiply by the opposite

OpenStudy (lgbasallote):

in this case, you have subtraction...so you multiply by addition

OpenStudy (anonymous):

Then my answer would be 40✔5

OpenStudy (lgbasallote):

close...but no

OpenStudy (lgbasallote):

if you multiply those... \[\implies \frac{5(8 + \sqrt 5)}{(8-\sqrt 5)(8+\sqrt 5)}\] expand the numerator by distribution...expand the denominator by difference of two squares

OpenStudy (lgbasallote):

anyway, i have to go now so i hope you luck in solving this

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

(8+√5) (8−√5 ) = ......?